# Towards AI Search Paradigm

### v1

@@ 3.2 Task Universe and the MCP Abstraction
Early tool-augmented LLM systems relied on vendor-specific “function-calling” JSON schemas such as those introduced by OpenAI (OpenAI, 2025a, b) and rapidly copied by many frameworks (LangChain, 2024). While simple, these ad-hoc contracts were coupled to one provider, lacked machine-typed guarantees, and made it impossible for independent agents to share tools or reason about cost, latency, or security across organizational boundaries. For AI Search whose planner must orchestrate heterogeneous knowledge look-ups, computations, and transformations in a single multi-step plan, such fragmentation is a major bottleneck.
Model-Context Protocol. MCP addresses this fragmentation by specifying a vendor-neutral, HTTP+JSON-RPC interface through which Servers expose tools and data, while Clients (i.e., LLMs or agents) discover, invoke, and monitor those tools in a secure, typed fashion (Anthropic, 2024; Schmid, 2025). Specifically, the protocol bundles four ingredients: (1) a manifest that advertises each endpoint’s name, semantic role, cost, and latency bounds; (2) machine-readable input/output schemas that ground the LLM’s function-calling tokens; (3) a capability handshake for tool discovery, and (4) an execution contract that guarantees idempotent, auditable calls.
@@ 3.3.1 Refinement of Tool API Documents
Considering the critical importance of precise tool description within AI Search paradigm, employing tool learning techniques to refine these documents represents an effective approach (Mialon et al., 2023; Qin et al., 2023b). Specifically, tool learning has emerged as a powerful approach for enhancing the ability of LLMs to tackle complex, real-world tasks (Schick et al., 2024; Qu et al., 2025a). By interacting with tools such as calculators, search engines, APIs, and web environments, LLMs could overcome key limitations arising from their fixed pretraining data and inherently passive text-in/text-out interface (Nakano et al., 2021; Qin et al., 2023a; M. Bran et al., 2024). This integration allows LLMs to acquire real-time information, perform dynamic actions, and extend their utility beyond traditional language modeling (Zhuang et al., 2024; Wang et al., 2024). To facilitate such interactions, LLMs are typically equipped with tool documentation that serves as contextual input, describing each tool’s functionalities, usage methods, and applicable scenarios (Shen et al., 2024; Song et al., 2023; Xu et al., 2023). The effectiveness of LLMs in leveraging external tools largely depends on the clarity, structure, and accuracy of the accompanying API documentation. While traditional tool documents are generally written for human developers, they often fall short of meeting the specific needs of LLMs by containing ambiguities, redundancies, or missing information, thereby impeding the accurate interpretation of tool functionalities and task execution (Chemero, 2023; Yuan et al., 2024). Given these limitations, a key question arises: How can we construct tool documentation that is better aligned with the needs of the Planner?
@@
To enhance the quality of tool documentation, the AI search utilizes an iterative refinement method DRAFT, which leverages the interactions between LLMs and external tools, as well as the feedback generated during these interactions, to progressively optimize tool documents (Qu et al., 2025b). As illustrated in Fig. 4, DRAFT involves three iterative phases: Experience Gathering, Learning from Experience, and Documentation Rewriting. Specifically, DRAFT first systematically simulates diverse use-cases, encompassing typical interactions, edge cases, error scenarios, and parameter limits. Such exhaustive exploratory interactions uncover gaps and inaccuracies in existing tool descriptions. Subsequently, DRAFT analyzes collected interaction data to identify discrepancies and ambiguities. Insights derived from practical tool interactions ensure that refinements align closely with actual usage contexts. This analytical process yields targeted recommendations aimed at correcting inaccuracies, clarifying ambiguities, and eliminating redundancy. Eventually, DRAFT integrates these recommendations, synthesizing refined tool descriptions specifically optimized for effective interpretation by LLMs. The method incorporates strategies that promote diverse exploratory coverage and adaptive termination criteria to prevent excessive iterations, ensuring concise yet comprehensive documentation. Next, we will describe the algorithm details of our proposed method as follows.
@@
where $\phi$ is a predefined similarity threshold. If the constraint is violated, the explorer discards the instance and reflects on the redundancy (Shinn et al., 2024), iteratively refining the generated query until a sufficiently novel query is generated.
@@
Recognizing that tool complexity varies, we adopt a tool-adaptive termination mechanism to decide when to stop iterations. We measure the change $\Delta$ between consecutive versions using a combination of the word-match metric BLEU score (Papineni et al., 2002) and the semantic-match metric cosine similarity, formulated as
@@ 3.3.2 Tool Clustering in MCP
To organize the tool embeddings into functionally coherent groups, the AI search system uses the $k$-means++ algorithm (Arthur and Vassilvitskii, 2006), which partitions the embedding space into $k$ distinct clusters. The optimization objective of the clustering process can be formulated as
@@ 3.3.3 Query-Oriented Tool Retrieval
However, the above LLM-based tool retrieval strategies, typically focusing on semantic matching between queries and individual tools (Xiong et al., 2020; Hofstätter et al., 2021; Gao and Callan, 2021; Izacard et al., 2021; Ma et al., 2022), inadequately address complex queries that necessitate collaborative tool use, resulting in incomplete or suboptimal outcomes. For example, a query like calculating the value of 5 ounces of gold plus 1 million AMZN stocks in CNY requires tools for gold pricing, stock valuation, and currency conversion. If any are missing, the answer is incomplete. Traditional dense retrieval methods, focused on semantic similarity, often retrieve redundant tools (e.g., multiple stock tools) while omitting others, failing to capture the need for tool collaboration. This underscores the often-overlooked importance of completeness in tool retrieval. Considering these factors, a central challenge arises: how can we design a retriever that selects a complete and task-relevant set of tools to support the Planner effectively?
To address the above challenge, the AI search system leverages a PLM-enhanced retrieval method COLT, which integrates the semantic and collaborative dimensions of tool functionality (Qu et al., 2024). Initially, semantic relationships between queries and tools are established through representation learning techniques based on pre-trained language models. Subsequently, COLT employs graph-based learning methods to capture and refine collaborative interactions among tools, queries, and task scenarios. By adopting dual-view graph structures that explicitly model these interactions, COLT effectively captures the high-order collaborative information between tools. Furthermore, to promote balanced retrieval across all relevant tools, the learning objective incorporates a list-wise multi-label loss, preventing over-reliance on individual tools. Next, we detail the proposed tool retrieval method as follows.
Sematic Learning. COLT first focuses on semantic representation learning through dense retrieval, a widely adopted approach in modern retrieval systems (Zhao et al., 2023a; Guo et al., 2022). It employs a bi-encoder architecture powered by PLMs, such as BERT (Kenton and Toutanova, 2019), to independently encode both the query $q$ and each tool description $t$ into dense vector embeddings. These embeddings, obtained via mean pooling over the final layer outputs, are then compared using cosine similarity to estimate their initial semantic relevance. The training process is performed using the InfoNCE loss (Gutmann and Hyvärinen, 2010; Xiong et al., 2020), which contrasts positive tool-query pairs against negatives:
$-\log\frac{e^{\mathrm{sim}(q,t^{+})}}{e^{\mathrm{sim}(q,t^{+})}+\sum_{j=1}^{k}e^{\mathrm{sim}(q,t_{j}^{-})}},$ (9)
@@ 3.4 DAG-based Task Planning
Specifically, the AI search system leverages a candidate tool API set $\mathcal{T}$ and the user query $q$ as contextual input of the Planner, in a single inference pass, and generates a task graph $G=(V,E)$, where $V$ is the set of vertices (i.e., sub-tasks) and $E$ refers to the set of dependency edges. Each vertex $v\in V$ represents an atomic and schedulable sub-task, which may either bind to an external tool or be fulfilled by the Executor using local LLM computation. For uniform specification, every vertex is equipped with the interface signature described as $\bigl{\langle}\text{arg}(v),\ \text{ret}(v)\bigr{\rangle}$, where $\text{arg}(v)$ declares the parameter slots that must be instantiated before execution of the sub-task, and $\text{ret}(v)$ describes the structure of the data returned afterwards. If the sub-task requires an external tool, the binding is recorded as $\tau(v)\in\mathcal{T}$; otherwise, $\tau(v)=\varnothing$ and the computation is performed locally.
@@
  • Atomic Sub-Tasks. For every vertex $v$ at most one tool type $\tau(v)$ can be bound, although that tool may be invoked multiple times inside the sub-task.
  • Sequential Dependency. Second, the transitive closure of $G$ must capture all sequential dependencies inherent in the semantic relations required to address $q$, thereby ensuring completeness while minimizing inter-task coupling.
To elicit a well-structured DAG $G$ in a single forward pass, the AI search system employs a chain-of-thought $\rightarrow$ structured-sketch prompting strategy. Concretely, the Planner is first encouraged to decompose $q$ implicitly through reasoning, annotating latent variables and dependencies. Then, the resulting draft of the reasoning is reorganized into a JSON standard format for DAG via instruction prompting. Unlike iterative ReAct-style reasoning, this approach produces a machine-readable, verifiable global plan simultaneously, substantially reducing token consumption and simplifying downstream parsing.
@@
In this way, DAG-based task planning externalizes multi-step reasoning into a structured task graph, achieving minimal context length, layer-wise parallelism, and full traceability. This design provides a principled, modular backbone for subsequent reflection mechanisms and reinforcement-learning optimization, which will be discussed in the remaining sections.
@@ 3.6 Planner Optimization with RL Strategy
    $\mathcal{R}_{Feedback}=\begin{cases}\alpha_{Feedback},&\text{if users accept the sub-tasks;}\\ \beta_{Feedback},&\text{if users reject the sub-tasks,}\end{cases}$ (12)
@@
    $\mathcal{R}_{Format}=\begin{cases}\alpha_{Format},&\text{if output is correctly formatted;}\\ \beta_{Format},&\text{if output is incorrectly formatted,}\end{cases}$ (13)
@@
Given the above rewards, the AI search system leverages Group Relative Policy Optimization (GRPO) (Shao et al., 2024) to optimize the Planner. Specifically, for a given complex query $q$, the AI search system samples a set of outputs $o=\{o_{1},o_{2},\cdots,o_{N}\}$ from the previous planner policy model $\pi_{\theta_{\text{old}}}$. Corresponding to these outputs, it computes a set of rewards $r=\{r_{1},r_{2},\cdots,r_{N}\}$ based on the rule-defined reward function (Eq. 15). The advantage for each token is defined as the normalized reward as
@@ 4.1 Background
Pointwise. These approaches utilize LLMs to independently assess the relevance of each candidate document with the given query. Formally, given a query $q$ and a retrieved document $\mathcal{D}_{i}$, we aim to instruct the LLM to generate a quantitative, comparable score for each query and candidate document pair, denoted as the basic labeling unit as $u(q,\mathcal{D}_{i})$. This score can be a graded label commonly used in human annotation or a logit value output by the model. Then, we apply the labeling unit to all the candidate documents and obtain the final rankings based on the sorted scores. For instance, some works directly prompt LLMs to generate whether the provided candidate document is relevant to the query and the documents are labeled based on the normalized likelihood of generating a “yes” response (Nogueira et al., 2020; Liang et al., 2022; Sun et al., 2023a). The other method involves employing LLMs to generate a relevant query for each candidate document and then re-rank these documents based on the likelihood of generating the actual query (Sachan et al., 2022). Another approach involves appending an output layer to the LLM’s representation and then employing SFT with human-annotated samples for re-ranking tasks. This process produces a scalar score that reflects the relative query-document relevance (Ma et al., 2024).
Pairwise. These approaches use LLMs to compare the relevance of a pair of candidate documents with the given query. Formally, given a query $q$ and a pair of retrieved documents $(\mathcal{D}_{i},\mathcal{D}_{j})$, we aim to ask the LLM to determine which document is more relevant to the given query $q$, denoted as the basic labeling unit $u(q,\mathcal{D}_{i},\mathcal{D}_{j})$. There are two principal approaches leveraging the pairwise score unit: the all-pairs method and the sorting-based method (Pradeep et al., 2021; Qin et al., 2023c). The all-pairs method generates all possible permutations of document pairs from the candidate documents set. Then, each pair is evaluated independently by the LLM, and a score function is then used to assign a score to each document based on the pairwise comparisons. The final ranking is determined by the total score each document receives. In contrast, the sorting-based method utilizes sorting algorithms, such as heap sort and bubble sort, to leverage efficient data structures for selectively comparing document pairs. This enables the most relevant documents to be quickly identified and placed at the top of the final ranking without the need for exhaustive comparison of all possible pairs. More advanced research, such as PRP-Graph (Luo et al., 2024), draws inspiration from real-world tournaments and incorporates graph structures into pairwise ranking to further improve efficiency.
@@
Listwise. These approaches utilize LLMs to directly rank multiple documents within a single window based on their relevance to the given query (Sun et al., 2023b; Liu et al., 2024; Chen et al., 2025a). Formally, given a query $q$ and a list of retrieved documents $\{\mathcal{D}_{1},\dots,\mathcal{D}_{m}\}$, we aim to ask the LLM to directly generate an ordered doc-list based on the relevance to the given query $q$, denoted as the basic labeling unit $u(q,\{\mathcal{D}_{1},\dots,\mathcal{D}_{m}\})$. Since the Listwise method achieves a good trade-off between effectiveness and efficiency for ranking multiple documents at one time, they are considered as the preferred prompting strategy for LLM ranking. In the following sections, we will detail two Listwise-based ranking methods in practice: RankGPT (Sun et al., 2023b) and TourRank (Chen et al., 2025a).

### v2

@@ 3.2 Task Universe and the MCP Abstraction
Early tool-augmented LLM systems relied on vendor-specific “function-calling” JSON schemas such as those introduced by OpenAI [openai2025responses, openaifunctioncalling] and rapidly copied by many frameworks [langchainagentprotocol]. While simple, these ad-hoc contracts were coupled to one provider, lacked machine-typed guarantees, and made it impossible for independent agents to share tools or reason about cost, latency, or security across organizational boundaries. For AI Search whose planner must orchestrate heterogeneous knowledge look-ups, computations, and transformations in a single multi-step plan, such fragmentation is a major bottleneck.
Model-Context Protocol. MCP addresses this fragmentation by specifying a vendor-neutral, HTTP+JSON-RPC interface through which Servers expose tools and data, while Clients (i.e., LLMs or agents) discover, invoke, and monitor those tools in a secure, typed fashion [anthropic2024mcp, schmid2025overview]. Specifically, the protocol bundles four ingredients: (1) a manifest that advertises each endpoint’s name, semantic role, cost, and latency bounds; (2) machine-readable input/output schemas that ground the LLM’s function-calling tokens; (3) a capability handshake for tool discovery, and (4) an execution contract that guarantees idempotent, auditable calls.
@@ 3.3.1 Refinement of Tool API Documents
Considering the critical importance of precise tool description within AI Search paradigm, employing tool learning techniques to refine these documents represents an effective approach [mialon2023augmented, qin2023tool]. Specifically, tool learning has emerged as a powerful approach for enhancing the ability of LLMs to tackle complex, real-world tasks [schick2024toolformer, qu2025tool]. By interacting with tools such as calculators, search engines, APIs, and web environments, LLMs could overcome key limitations arising from their fixed pretraining data and inherently passive text-in/text-out interface [nakano2021webgpt, qin2023webcpm, m2024augmenting]. This integration allows LLMs to acquire real-time information, perform dynamic actions, and extend their utility beyond traditional language modeling [zhuang2023toolchain, wang2024llms]. To facilitate such interactions, LLMs are typically equipped with tool documentation that serves as contextual input, describing each tool’s functionalities, usage methods, and applicable scenarios [shen2024hugginggpt, song2023restgpt, xu2023tool]. The effectiveness of LLMs in leveraging external tools largely depends on the clarity, structure, and accuracy of the accompanying API documentation. While traditional tool documents are generally written for human developers, they often fall short of meeting the specific needs of LLMs by containing ambiguities, redundancies, or missing information, thereby impeding the accurate interpretation of tool functionalities and task execution [chemero2023llms, yuan2024easytool]. Given these limitations, a key question arises: How can we construct tool documentation that is better aligned with the needs of the Planner?
@@
To enhance the quality of tool documentation, the AI search utilizes an iterative refinement method DRAFT, which leverages the interactions between LLMs and external tools, as well as the feedback generated during these interactions, to progressively optimize tool documents [quexploration]. As illustrated in Fig. 4, DRAFT involves three iterative phases: Experience Gathering, Learning from Experience, and Documentation Rewriting. Specifically, DRAFT first systematically simulates diverse use-cases, encompassing typical interactions, edge cases, error scenarios, and parameter limits. Such exhaustive exploratory interactions uncover gaps and inaccuracies in existing tool descriptions. Subsequently, DRAFT analyzes collected interaction data to identify discrepancies and ambiguities. Insights derived from practical tool interactions ensure that refinements align closely with actual usage contexts. This analytical process yields targeted recommendations aimed at correcting inaccuracies, clarifying ambiguities, and eliminating redundancy. Eventually, DRAFT integrates these recommendations, synthesizing refined tool descriptions specifically optimized for effective interpretation by LLMs. The method incorporates strategies that promote diverse exploratory coverage and adaptive termination criteria to prevent excessive iterations, ensuring concise yet comprehensive documentation. Next, we will describe the algorithm details of our proposed method as follows.
@@
where $\phi$ is a predefined similarity threshold. If the constraint is violated, the explorer discards the instance and reflects on the redundancy [shinn2024reflexion], iteratively refining the generated query until a sufficiently novel query is generated.
@@
Recognizing that tool complexity varies, we adopt a tool-adaptive termination mechanism to decide when to stop iterations. We measure the change $\Delta$ between consecutive versions using a combination of the word-match metric BLEU score [papineni-etal-2002-bleu] and the semantic-match metric cosine similarity, formulated as
@@ 3.3.2 Tool Clustering in MCP
To organize the tool embeddings into functionally coherent groups, the AI search system uses the $k$-means++ algorithm [arthur2006k], which partitions the embedding space into $k$ distinct clusters. The optimization objective of the clustering process can be formulated as
@@ 3.3.3 Query-Oriented Tool Retrieval
However, the above LLM-based tool retrieval strategies, typically focusing on semantic matching between queries and individual tools [xiong2020approximate, hofstatter2021efficiently, gao2021unsupervised, izacard2021unsupervised, ma2022pre], inadequately address complex queries that necessitate collaborative tool use, resulting in incomplete or suboptimal outcomes. For example, a query like calculating the value of 5 ounces of gold plus 1 million AMZN stocks in CNY requires tools for gold pricing, stock valuation, and currency conversion. If any are missing, the answer is incomplete. Traditional dense retrieval methods, focused on semantic similarity, often retrieve redundant tools (e.g., multiple stock tools) while omitting others, failing to capture the need for tool collaboration. This underscores the often-overlooked importance of completeness in tool retrieval. Considering these factors, a central challenge arises: how can we design a retriever that selects a complete and task-relevant set of tools to support the Planner effectively?
To address the above challenge, the AI search system leverages a PLM-enhanced retrieval method COLT, which integrates the semantic and collaborative dimensions of tool functionality [qu2024towards]. Initially, semantic relationships between queries and tools are established through representation learning techniques based on pre-trained language models. Subsequently, COLT employs graph-based learning methods to capture and refine collaborative interactions among tools, queries, and task scenarios. By adopting dual-view graph structures that explicitly model these interactions, COLT effectively captures the high-order collaborative information between tools. Furthermore, to promote balanced retrieval across all relevant tools, the learning objective incorporates a list-wise multi-label loss, preventing over-reliance on individual tools. Next, we detail the proposed tool retrieval method as follows.
Sematic Learning. COLT first focuses on semantic representation learning through dense retrieval, a widely adopted approach in modern retrieval systems [zhao2022dense, guo2022semantic]. It employs a bi-encoder architecture powered by PLMs, such as BERT [kenton2019bert], to independently encode both the query $q$ and each tool description $t$ into dense vector embeddings. These embeddings, obtained via mean pooling over the final layer outputs, are then compared using cosine similarity to estimate their initial semantic relevance. The training process is performed using the InfoNCE loss [gutmann2010noise, xiong2020approximate], which contrasts positive tool-query pairs against negatives:
$\log\frac{e^{\mathrm{sim}(q,t^{+})}}{e^{\mathrm{sim}(q,t^{+})}+\sum_{j=1}^{k}e^{\mathrm{sim}(q,t_{j}^{-})}},$ (9)
@@ 3.4 DAG-based Task Planning
Specifically, the AI search system leverages a candidate tool API set $\mathcal{T}$ and the user query $q$ as contextual input of the Planner, in a single inference pass, and generates a task graph $G=(V,E)$, where $V$ is the set of vertices (i.e., sub‑tasks) and $E$ refers to the set of dependency edges. Each vertex $v\in V$ represents an atomic and schedulable sub-task, which may either bind to an external tool or be fulfilled by the Executor using local LLM computation. For uniform specification, every vertex is equipped with the interface signature described as $\bigl\langle\text{arg}(v),\ \text{ret}(v)\bigr\rangle$, where $\text{arg}(v)$ declares the parameter slots that must be instantiated before execution of the sub-task, and $\text{ret}(v)$ describes the structure of the data returned afterwards. If the sub‑task requires an external tool, the binding is recorded as $\tau(v)\in\mathcal{T}$; otherwise, $\tau(v)=\varnothing$ and the computation is performed locally.
@@
  • Atomic Sub-Tasks. For every vertex $v$ at most one tool type $\tau(v)$ can be bound, although that tool may be invoked multiple times inside the sub‑task.
  • Sequential Dependency. Second, the transitive closure of $G$ must capture all sequential dependencies inherent in the semantic relations required to address $q$, thereby ensuring completeness while minimizing inter‑task coupling.
To elicit a well‑structured DAG $G$ in a single forward pass, the AI search system employs a chain‑of‑thought $\rightarrow$ structured‑sketch prompting strategy. Concretely, the Planner is first encouraged to decompose $q$ implicitly through reasoning, annotating latent variables and dependencies. Then, the resulting draft of the reasoning is reorganized into a JSON standard format for DAG via instruction prompting. Unlike iterative ReAct-style reasoning, this approach produces a machine-readable, verifiable global plan simultaneously, substantially reducing token consumption and simplifying downstream parsing.
@@
In this way, DAG-based task planning externalizes multi-step reasoning into a structured task graph, achieving minimal context length, layer-wise parallelism, and full traceability. This design provides a principled, modular backbone for subsequent reflection mechanisms and reinforcement‑learning optimization, which will be discussed in the remaining sections.
@@ 3.6 Planner Optimization with RL Strategy
[…]
@@
[…]
@@
Given the above rewards, the AI search system leverages Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical] to optimize the Planner. Specifically, for a given complex query $q$, the AI search system samples a set of outputs $o=\{o_{1},o_{2},\cdots,o_{N}\}$ from the previous planner policy model $\pi_{\theta_{\text{old}}}$. Corresponding to these outputs, it computes a set of rewards $r=\{r_{1},r_{2},\cdots,r_{N}\}$ based on the rule-defined reward function (Eq. 15). The advantage for each token is defined as the normalized reward as
@@ 4.1 Background
Pointwise. These approaches utilize LLMs to independently assess the relevance of each candidate document with the given query. Formally, given a query $q$ and a retrieved document $\mathcal{D}_{i}$, we aim to instruct the LLM to generate a quantitative, comparable score for each query and candidate document pair, denoted as the basic labeling unit as $u(q,\mathcal{D}_{i})$. This score can be a graded label commonly used in human annotation or a logit value output by the model. Then, we apply the labeling unit to all the candidate documents and obtain the final rankings based on the sorted scores. For instance, some works directly prompt LLMs to generate whether the provided candidate document is relevant to the query and the documents are labeled based on the normalized likelihood of generating a “yes” response [nogueira2020document, Liang2022HolisticEO, sun2023instruction]. The other method involves employing LLMs to generate a relevant query for each candidate document and then re-rank these documents based on the likelihood of generating the actual query [Sachan2022ImprovingPR]. Another approach involves appending an output layer to the LLM’s representation and then employing SFT with human-annotated samples for re-ranking tasks. This process produces a scalar score that reflects the relative query-document relevance [10.1145/3626772.3657951].
Pairwise. These approaches use LLMs to compare the relevance of a pair of candidate documents with the given query. Formally, given a query $q$ and a pair of retrieved documents $(\mathcal{D}_{i},\mathcal{D}_{j})$, we aim to ask the LLM to determine which document is more relevant to the given query $q$, denoted as the basic labeling unit $u(q,\mathcal{D}_{i},\mathcal{D}_{j})$. There are two principal approaches leveraging the pairwise score unit: the all-pairs method and the sorting-based method [Pradeep2021TheED, Qin2023LargeLM]. The all-pairs method generates all possible permutations of document pairs from the candidate documents set. Then, each pair is evaluated independently by the LLM, and a score function is then used to assign a score to each document based on the pairwise comparisons. The final ranking is determined by the total score each document receives. In contrast, the sorting-based method utilizes sorting algorithms, such as heap sort and bubble sort, to leverage efficient data structures for selectively comparing document pairs. This enables the most relevant documents to be quickly identified and placed at the top of the final ranking without the need for exhaustive comparison of all possible pairs. More advanced research, such as PRP-Graph [Luo2024PRPGraphPR], draws inspiration from real-world tournaments and incorporates graph structures into pairwise ranking to further improve efficiency.
@@
Listwise. These approaches utilize LLMs to directly rank multiple documents within a single window based on their relevance to the given query [Sun2023IsCG, liu2024sliding, Chen2024TourRankUL]. Formally, given a query $q$ and a list of retrieved documents $\{\mathcal{D}_{1},\dots,\mathcal{D}_{m}\}$, we aim to ask the LLM to directly generate an ordered doc-list based on the relevance to the given query $q$, denoted as the basic labeling unit $u(q,\{\mathcal{D}_{1},\dots,\mathcal{D}_{m}\})$. Since the Listwise method achieves a good trade-off between effectiveness and efficiency for ranking multiple documents at one time, they are considered as the preferred prompting strategy for LLM ranking. In the following sections, we will detail two Listwise-based ranking methods in practice: RankGPT [Sun2023IsCG] and TourRank [Chen2024TourRankUL].